# A Computational Theory and Semi-Supervised Algorithm for Clustering

### v1

@@ 2.2 Computational Theory
The clustering problem is generally understood to be the grouping of portions of data that are similar in a particular way, but separate from other such groups. The process has important applications in data mining where the output can be used—amongst other purposes—for segmentation, summarisation and classification. The input to the problem is assumed to be in a dataframe format, where examples are represented as rows and feature values as columns of numbers. The desired output is a one-dimensional array of labels that identify which cluster each example belongs to and in addition there may be a membership score. Note that a specific cluster label is required for the unknown or anomalous group. Unfortunately, there is not enough information available in the input to get a unique desired output directly and hence any partitioning of the data could be returned. The problem is ill-posed, and we require additional external information to constrain the possible solutions; we derive this from the human experience of grouping data points in low dimensisonal spaces.
The process of clustering is distinct from partitioning; subsets of data points may be grouped into a cluster but not necessarily all points should be assigned to a cluster. This then provides our first natural and most important constraint that similar points should be clustered together, but anomalies with respect to every grouping and measure, must not belong to any cluster. The second assumption is that an object that is part of a cluster has both a membership and degree of membership. In other words even within a single group all observations have a level of belonging. For example, points that are further away from the centroid or occupy less dense regions of space than others.
@@
This definition needs to be made precise and requires some additional explanation and support. Assuming that the realworld physically imposes natural groupings of interest, the term anomaly is taken to be that defined by Mohammad [8].
@@
Anomalies, according to this definition can be computed precisely, and an algorithm under a particular representation of Euclidean distance from the median is provided by Mohammad [8]. This not only provides us with a classification but also a score defining the degree of anomalousness. Hence, clusters being simply the complement of the set of anomalies in the data are also computable and have an associated score. This also satisfies a third constraint that is taken as an assumption: what is computed must have parameters or thresholds that adapt to the data, and are not fixed unless they are universally useable. This is reasonable to assume since data sets, and indeed clusters within data sets, can vary in terms of numbers, size, density or shape and an algorithm needs to be able to best adapt to such differences. However, the constraint is unlikely to be universally satisfied across all data sets due to the range of possible data distributions and the specificity of each adaptation.
The clustering process is approached from the perspective of removing anomalies as it provides a more intuitive and appreciable framework. This viewpoint allows for easier understanding and handling of the subject and algorithms, emphasising that points classified as not belonging to any group are considered anomalies.
Gestalt groupings play an important role since clustering is often described as finding natural groupings in the data. (Figure 1 illustrates some examples of ‘natural’ clusters by proximity.) Thus, it can be argued that satisfactory clustering results are not completely aribtrary, in that though a staggeringly large number of possible cluster outputs can be returned, it is only a relative few that are acceptable. Thus, the groupings that we care about and consider natural are limited and are conjectured to include grouping by the gestalt laws such as proximity, good continuity and similarity. These groupings should agree with what the common man sees and believes to be true. Clustering is in the eye of the beholder is an old adage, but the eye being human is constrained by only perceiving certain patterns that correspond to prior grouping laws of interest.
@@ 2.4 Algorithm
Given a representation of the data points—which is defined as the Euclidean distance from the associated median—and a small subsample of labelled data, this section presents an algorithm for transforming the input into the desired output. The primary computational task involves detecting and removing anomalies from a set of data points, while simultaneously expanding clusters by incorporating non-anomalous points. Thus, anomaly detection lies at the core of the philosophy and method employed.
@@
Here, we also acknowledge an additional constraint arising from the human ability to quickly perceive clusters in low-dimensional data, as well as the requirement to efficiently compute clusters over large realworld datasets. Namely, that the developed algorithm must be fast to compute and scale efficiently with increasing numbers of examples and dimensions of the data. Many clustering algorithms that yield satisfactory results with small datasets or low-dimensional data often struggle to meet these requirements, resulting in unacceptably long computation times.
To begin with, the clustering process requires users to provide a small subsample of known cluster assignments, possibly obtained randomly, that are largely accurate. These assignments serve as the initial seeds and guides for the clustering, while all other points are initially considered anomalies. Subsequently, the algorithm iterates over the clusters in order of the tightest concentration of points. For each cluster, anomalies are identified and ejected from the cluster, being reassigned to the anomaly category. Following this step, the algorithm examines all the anomalous points to determine if they should be considered anomalies or potentially belong to the cluster. Points that are determined to be non-anomalous have their label assignments changed to match the current cluster, while the labels of anomalies remain unchanged. This process of ejecting anomalies and expanding clusters continues iteratively until no further changes in the clustering occur or a maximum number of iterations have completed. The algorithm’s pseudocode is given below where Mohammad’s anomaly detection algorithm, called Perception, is used to detect anomalies parameter-free:
Input : Data points, subsample of labeled data, $max\_n\_iterations=1000$
Output : Cluster labels and membership score
@@
1Initialise all unlabelled points as anomalies with label $-1$;
2 Order the labelled cluster groups by minimum sum of squared error;
3 $counter=0$;
@@
6             $clf=\text{Perception()}$() ;
7             $clf.fit(\text{cluster})$;
8             $clf.predict(\text{cluster})$;
@@ 6             $clf=\text{Perception()}$() ;
10                   Remove anomalies from the cluster and assign them to the anomaly category;
11
@@ 10                   Remove anomalies from the cluster and assign them to the anomaly category;
15                   prediction, score = $clf.predict(\text{$anomalous\ point$})$;
@@
17                         Change point label assignment to match the current cluster;
18
@@ 17                         Change point label assignment to match the current cluster;
21                         Leave anomaly label unchanged;
22
@@ 28      $counter=counter+1$;
31Re-fit the Perception() model on each cluster and compute scores;
32 return labels and membership scores of all examples;
@@ 3 Experimental Results
This section provides the results of the clustering algorithm over several data sets. This includes one-dimensional and two-dimensional synthetic data, and then a selection of publicly available data sets. Although the presented method is semi-supervised, comparisons are not made against COPK-means or PCK-means. This is because their results were remarkably similar to standard $k$-means, but with longer execution times and the need for additional constraints. Instead, comparisons are made with $k$-means (with the given $k$ used where possible) and DBSCAN ($mps$ and $eps$ decided as in the guidelines of the original paper). The implementation of these algorithms is done using the scikit-learn library [9]. These algorithms were chosen due to their widespread appeal and usage when users encounter clustering problems, aiming to demonstrate the effectiveness of our proposed method.
Evaluation of clustering algorithms is notoriously difficult because of the large number of measures from which to choose, and because certain measures favour certain algorithms due to the nature of the optimisation problem being solved. Partioning algorithms will also give very different results to clustering algorithms that report anomalies (such as the one presented in this paper), and may also appear artificially superior. This is because many data sets assign all examples to some cluster, and do not highlight those that are actually anomalies. Hence, even the ground truth labelling may not always be an accurate reflection of reality; for example in the generation of gaussian clusters it can be argued that there are outlying points with respect to the central mass of points of each cluster (e.g., Figure 9). In addition, it is important to consider that a dataset can be clustered in multiple ways or based on different features. Therefore, even with the availability of labels for evaluation, a clustering method may produce valid and effective clusters but which do not align with the pre-existing labels. Consequently, this work adopts a qualitative approach to address the clustering evaluation problem.
@@ 3.1 One-dimensional Synthetic Data
The first data set presented is composed of three gaussian distributed clusters, each of size $10000$ examples with centers $[0,50,100]$ and standard deviations of $[1,3,6]$. Added to this set are a few additional isolated local anomalies, a few mislabelled examples, and a small globally anomalous cluster. Figure 1 illustrates the data distribution where each of the three clusters are labelled and coloured—black, blue and green—and anomalies shown in red. Note however that anomalies resulting from the generating gaussian distribution for each of the clusters are not highlighted as such since it would require selecting an anomaly detection algorithm which may give results different to another, and that such algorithms often require appropriate but unknown parameters. In addition to the labelled scatter plot, a histogram is also provided to show the relative densities which is otherwise hidden in the plot.
@@
Figure 1: Three labeled gaussian distributed clusters with a few additional mislabelled examples, some isolated anomalies and a small anomalous cluster.
@@
A tiny fraction of the data—approximately $150$ examples ($0.5\%$) together with their labels—are randomly chosen from the data (excluding the small anomalous cluster). These labels act as the seeds for the proposed semi-supervised clustering algorithm and are expected to have been provided by the user after familiarising themselves and interacting with the data and domain. Note that the seeds need not cover all clusters since unrecognised clusters will simply be reported as anomalies and can be analysed in subsequent rounds. The results of the algorithm are shown in Figure 2 where three clusters are identified along with the isolated anomalies and small anomalous cluster; the latter of which can be investigated by the user in order to decide if it satsifies her criteria of being a new cluster. Note that the algorithm importantly flags the points in the outer regions of each cluster as anomalous. Closer inspection of one of the clusters and its histogram distribution illustrates how this can be desirous since these points may be considered outliers compared to the main mass of the cluster (Figure 3). Thus, the proposed method carries out a clustering of data as opposed to partitioning all data points into some group. Indeed, $k$-means partitions all data into one of three clusters—including the anomalies (Figure 4), while DBSCAN ($mps=10,eps=0.114$) on the other hand detects many anomalies, but also returns many more clusters embedded within the three main groups (Figure 5).
Figure 2: Results of the proposed semi-supervised clustering algorithm. Note that the three clusters are found using the seed labels. Anomalies are also identified as being unexpectedly far from any of the clusters, and the anomalous cluster of points is also not forced into any labelled cluster.
@@
Figure 3: A close up view of cluster 1 from Figure 2 showing the identification of not only anomalies and mislabelled data, but also those that might be considered rare or unusual in a gaussian distributed cluster of points.
@@
Figure 5: DBSCAN results are shown (without the histogram for clarity). Note that it has a concept of anomalies and thus identifies many. However, in this particular example too many additional clusters are returned amongst the three main clusters.

### v2

@@ 2.2 Computational Theory
The clustering problem is generally understood to be the grouping of portions of data that are similar in a particular way, but separate from other such groups. The process has important applications in data mining where the output can be used—amongst other purposes—for segmentation, summarisation and classification. The input to the problem is assumed to be in a dataframe format, where examples are represented as rows and numerical feature values as columns. The desired output is a one-dimensional array of labels that identify which cluster each example belongs to and in addition there may be a membership score. Note that a specific cluster label is required for the unknown or anomalous group. Unfortunately, there is not enough information available in the input to get a unique desired output directly and hence any partitioning of the data could be returned. The problem is ill-posed, and we require additional external information to constrain the possible solutions; we derive this from the human experience of grouping data points in low dimensisonal spaces. Thus, a key research insight in clustering lies in identifying principled constraints and assumptions that meaningfully reflect the structure of the data and the objectives of the task.
The process of clustering is distinct from partitioning; subsets of data points may be grouped into a cluster but not necessarily all points should be assigned to a cluster. This then provides our first natural and most important constraint that similar points should be grouped together, but anomalies with respect to every grouping and measure, must not belong to any cluster. The second assumption is that an object that is part of a cluster has both a membership and degree of membership. In other words even within a single group all observations have a level of belonging, e.g., points that are further away from the centroid or occupy less dense regions of space than others.
@@
This definition needs to be made precise and requires some additional explanation and support in order for computation. The Gestalt groupings play an important role since clustering is often described as finding natural groupings in the data. (Figure 1 illustrates some examples of ‘natural’ clusters by proximity.) Thus, it can be argued that satisfactory clustering results are not completely aribtrary, in that though a staggeringly large number of possible cluster outputs can be returned, it is only a relative few that are acceptable. Thus, the groupings that we care about and consider natural are limited and are conjectured to include grouping by the gestalt laws such as proximity, good continuity and similarity. These groupings should agree with what the common man sees and believes to be true. Clustering is in the eye of the beholder is an old adage, but the eye being human is constrained by only perceiving certain patterns that correspond to prior grouping laws of interest. Assuming that the realworld physically imposes natural groupings of interest, the term anomaly is taken to be that defined by Mohammad [22].
@@
Anomalies, according to this definition can be computed precisely, and an algorithm under a particular representation of Euclidean distance from the median is provided by Mohammad [22]. This not only provides us with a classification but also a score defining the degree of anomalousness. Hence, clusters being simply the complement of the set of anomalies in the data are also computable and each example has an associated score. This also provides us another constraint: what is computed must have parameters or thresholds that adapt to the data, and are not fixed unless they are universally useable. This is reasonable to assume since data sets, and indeed clusters within data sets, can vary in terms of numbers, size, density or shape and an algorithm needs to be able to best handle and adapt to such differences. However, the constraint is unlikely to be universally satisfied across all data sets due to the range of possible data distributions and the specificity of each adaptation.
The clustering process is approached from the perspective of removing anomalies as opposed to forming groups, as it provides a more intuitive and appreciable framework. This viewpoint allows for simpler understanding and handling of the subject and algorithms, emphasising that points classified as not belonging to any group are considered anomalies.
@@ 2.4 Algorithm
Given a representation of the data points—which is defined as the Euclidean distance from the associated median—and a small subsample of labelled data, this section presents an algorithm for transforming the input into the desired output. The main computational task is to detect and remove anomalies from a set of data points, while simultaneously expanding clusters by incorporating non-anomalous points. Thus, anomaly detection lies at the core of the philosophy and method employed.
@@
Here, we also acknowledge an additional constraint arising from the human ability to quickly perceive clusters in low-dimensional data, as well as the requirement to efficiently compute clusters over large realworld datasets. Namely, that the developed algorithm must be fast to compute and scale efficiently with increasing numbers of examples and dimensions of the data. Many clustering algorithms that yield satisfactory results with small datasets or low-dimensional data often struggle to meet these requirements, resulting in unacceptably long computation times. Our approach prioritises computational efficiency to overcome these limitations.
The clustering method presented in this paper employs the perception anomaly detection algorithm due to several advantageous properties: it is parameter-free, adapts naturally to numerical data, is centroid-based for general applicability, and demonstrates computational efficiency on large datasets. Furthermore, it achieves state-of-the-art performance over metrics such as F1 score, AUC, precision, and recall on real-world data [22]. An explanation of the algorithm with a simple sketch is given as follows, while full details can be found in [22]. Consider a list of integers $V=[v_{1},v_{2},\dots,v_{W}]$ in which to detect anomalies, and assume they have been derived from a binary indicator stream $D$ composed of $1^{\prime}s$ and $0^{\prime}s$ where the stream was divided into $W$ fixed-length windows $w_{1},w_{2},\dots,w_{W}$, each of length $L$ (which can simply be assumed to be the largest integer in $V$). We can sum all the indicators in each window to have $V=[v_{1},v_{2},\dots,v_{W}]=[sum(w_{1}),sum(w_{2}),\dots,sum(w_{W})]$, and let $S=sum(V)$ be the sum of all the indicators across all the windows.
The a-contrario assumption and model is that all the indicators ($1^{\prime}s$ of total sum $S$) are uniformly and independently randomly distributed across the $W$ windows. Under this model, the quality of interest is the unexpectedly large number of indicators that appear in a window. This is formalised by computing the expected number of $n$-tuples $\mathbb{E}(C_{n})$ under the a-contrario distribution as a measure of how anomalous the number of elements in a particular window are. Thus, we are interested in events where the expected value is $<1$, which acts as a natural threshold for whether we expect the event to occur given the data parameters. This threshold is both principled and data-driven: if an event is unexpected under the assumed random distribution but appears in the data, it is, by definition, anomalous. Thus, the general formula for measuring the unexpectedness of each $n$-tuple and flagging an anomaly in $V$ is,
$\mathbb{E}(C_{n})={S\choose n}\frac{1}{W^{n-1}}<1$ (1)
This calculation gives an immediate binary output on the training data to class each window (value from V) as anomalous or benign. There are no parameters left to the user to specify. The use of expectations handles the probability thresholding problem to immediately arrive at the anomalous events, which can subsequently be ranked with smaller expectations corresponding to more meaningful (unexpected) events. The anomaly detection method easily extends to the multi-dimensional case, and the general training and test process together with actual implementation details is described in algorithms 1 and 2.
Input: $X$ input $l\times m$ dimensional array of numbers
$acc$ decimal place accuracy of input data (default 4)
$metric$ distance metric to be used (default Euclidean)
Output: $S$ sum of indicators
$W$ number of windows
$multiDimMed$ median values of each feature of X
$med$ median of one-dimensional data
$\mu$ mean of each feature of X
$\sigma$ standard deviation of feature of X
1
2if $dimension(X)>1$ then
3       $X_{s},\mu,\sigma\leftarrow standardise(X)$
4       $multiDimMed\leftarrow median(X_{s},axis=column)$
5       $X_{d}\leftarrow distance(X_{s},multiDimMed,metric)$
6       $X\leftarrow X_{d}$
7 end if
[…]
9 $med\leftarrow round(median(X_{g}))$
10 $X_{f}\leftarrow|X_{g}-med|$
11 $S\leftarrow sum(X_{f})$
12 $W\leftarrow length(X_{f})$
13
Algorithm 1 PerceptionTrain
1
Input: $X$ input $l\times m$ dimensional array of numbers to test
$S,W,multiDimMed,med,acc,metric,\mu,\sigma$ from training stage
Output: $Z$ array of prediction scores for each example row in $X$
$Y’$ ’ - binary array of predictions for $X$ (1: anomaly, 0: normal)
2
3if $dimension(X)>1$ then
4       $X_{s}\leftarrow standardise_{\mu,\sigma}(X)$
5       $X_{d}\leftarrow distance(X_{s},multiDimMed,metric)$
6       $X\leftarrow X_{d}$
7 end if
[…]
9 $X_{f}\leftarrow|X_{g}-med|$ Initialise $Y^{\prime}$ to zero array of length $l$
10 for $i=0$ to $l-1$ do
11       $n\leftarrow X_{f}[i]$
12       $Z[i]\leftarrow-1/S\times(log{S\choose n}-(n-1)\times log(W))$
13       if $Z[i]>0$ then
14             $Y’[i]\leftarrow 1$
15       end if
16
17 end for
Algorithm 2 PerceptionTest
The proposed clustering method begins with a small, user-provided subsample of data points with known and mostly accurate cluster assignments which may be drawn randomly from the dataset. These assignments act as seeds to initialise and guide the clustering process, while all remaining points are initially considered anomalies and assigned a label of $-1$. The clusters are then ordered by their compactness, as measured by the minimum sum of squared errors. For each cluster, the perception algorithm is applied to identify potential anomalies within it. Any detected anomalies are ejected from the cluster and reassigned the anomaly label $-1$. The perception algorithm is then re-fit to the updated cluster, and predictions are made over all current anomalies to determine whether any should be added to the cluster. This process over all clusters iterates until the cluster assignments stabilise (which typically occurs after only a few iterations) or a predefined iteration limit is reached. As a final step, the perception algorithm is re-applied to each cluster to compute a membership score for each assigned data point, quantifying its degree of belonging. The full procedure is detailed in Algorithm 3. (Note that application of the clustering method requires users to standardise the data to ensure features are on a similar scale. Any new data to test the method upon also requires the same standardisation parameters as those used during training.)
The clustering method iteratively assigns points to clusters by fitting and applying the perception anomaly detection method within each cluster. Since the perception has a computational complexity of $O(n\cdot d)$ for $n$ points and $d$ dimensions, running the perception method once per cluster per iteration results in a total cost of $O(n\cdot d)$ per iteration for the entire dataset. Over a maximum of $t$ iterations, which is typically small, this leads to an overall complexity of $O(t\cdot n\cdot d)$. Additionally, the final fitting and scoring step involves running perception on all clusters across the entire dataset, incurring an extra cost of $O(k\cdot n\cdot d)$, where $k$ is the number of clusters. Since usually $k\ll n$, the method scales approximately linearly with respect to the number of points and features, making it computationally efficient and suitable for moderate to large datasets.
The proposed clustering method offers a range of properties that meet our design goals while delivering practical utility. It integrates clustering with anomaly detection, allowing points that do not correspond to any known group to remain unassigned. These unlabelled points can be retained for exploratory analysis, potentially uncovering new or emerging clusters. By employing the Perception anomaly detector, the method adapts naturally to the structure of the data and requires no user-defined parameters, making it simple and robust in practice. It operates with minimal supervision, needing only a small subset of labelled examples to initiate the clustering process. For each assigned point, the method provides both a discrete cluster label and a membership score, indicating the strength of association and supporting nuanced downstream analysis. The approach also generalises well to unseen data: new points can be classified by evaluating their anomaly scores with respect to each learned cluster, enabling consistent and reliable prediction.
Input: $X$ Data points to cluster,
$X_{l}$ subsample of labelled data from $X$,
$max\_n\_iterations=1000$
Output: $L$ Cluster or anomaly label for each example in X
$M$ Cluster Membership score for each example in X
@@ 14             $Y’[i]\leftarrow 1$
1Initialise all unlabelled points ($X-X_{l}$) as anomalies with label $-1$
2 Initialise all labelled points ($X_{l}$) to a cluster number
3 Order the labelled cluster groups by minimum sum of squared error
4 $counter=0$
5
@@ 4 $counter=0$
8             $clf=\text{Perception()}$()
9             $clf.fit(\text{cluster})$
10             $clf.predict(\text{cluster})$
@@ 8             $clf=\text{Perception()}$()
12                   Remove anomalies from the cluster and assign them to the anomaly category
13
@@ 12                   Remove anomalies from the cluster and assign them to the anomaly category
17                   prediction, score = $clf.predict(\text{$anomalous\ point$})$
@@
19                         Change point label assignment to match the current cluster
20
@@ 19                         Change point label assignment to match the current cluster
23                         Leave anomaly label unchanged
24
@@ 30      $counter=counter+1$
33
34Re-fit the Perception() model on each cluster and compute scores
35 return labels $L$ and membership scores $M$ of all examples
@@ 3 Experimental Results
This section presents the evaluation results of the proposed clustering algorithm across a variety of datasets, including one-dimensional and two-dimensional synthetic datasets, as well as several publicly available benchmark datasets. We compare our method against a selection of representative clustering algorithms: $k$-means, DBSCAN, Agglomerative Clustering, Gaussian Mixture Models (GMMs), Seeded $k$-means (S-KM), Constrained $k$-means (C-KM), COP-KMeans (COPKM), and Deep Embedded Clustering (DEC). These methods were chosen because they are well-established baselines in the domains of unsupervised, semi-supervised, and deep clustering, and because reliable, well-written implementations are available. Implementation of $k$-means, Seeded $k$-means, DBSCAN, Agglomerative Clustering, and GMMs were obtained or adapted from the scikit-learn library [23]. For Constrained $k$-means and COP-KMeans, we used open-source implementations from the Python packages described in [17, 1]. The DEC model was implemented using the publicly available ClustPy package [15]. For all methods that require the number of ground truth clusters as input, we provided it explicitly. Although this could be seen as introducing a semi-supervised element, our intention was to ensure that algorithms like $k$-means were given their best chance to perform optimally with respect to this critical parameter. All other algorithm parameters were left at their default settings, reflecting the challenges of parameter tuning and the sensitivity of some methods to such choices in the absence of labelled data. The only exception here was for DEC, where the number of pretrain and clustering epochs were set to $10$ due to high computation time even at these low settings.
Evaluating clustering algorithms is inherently challenging due to the wide range of available metrics and the fact that many of these measures inherently favour certain types of algorithms, depending on the optimisation criteria they employ. For example, partitioning algorithms often produce results that differ significantly from those of clustering methods designed to detect anomalies—such as the approach proposed in this paper—and may appear to outperform them under conventional evaluation metrics. This is partly because many benchmark datasets assign all instances to a cluster, without explicitly labelling anomalous or outlying points. As a result, even the so-called ground truth may not accurately reflect the underlying structure of the data. For instance, in synthetic Gaussian mixtures, it is reasonable to consider some points as outliers relative to the core density of each cluster (see Figure 9). Furthermore, it is important to recognise that a dataset can often be clustered in multiple meaningful ways, depending on the features or criteria used. Consequently, even when ground truth labels are available, a clustering algorithm may discover valid and insightful groupings that do not align with the predefined labels. It has thus been recognised that human (with domain expertise) evaluation would provide a more meaningful measure of performance; however, this is a difficult time and resource consuming method. In light of these considerations, this work first adopts a qualitative evaluation strategy to better understand the patterns uncovered by the algorithms, followed by a quantitative automatic analysis using established external validation metrics to benchmark their performance.
@@ 3.1 One-dimensional Synthetic Data
The first dataset, named 1d_gauss, consists of three Gaussian-distributed clusters, each of sizes $[10000,5000,2500]$ examples with centers $[0,50,100]$ and standard deviations of $[1,3,8]$. Added to this set are a few additional isolated local anomalies, a few mislabelled examples, and a small globally anomalous cluster. Figure 1 illustrates the data distribution where each of the three clusters and anomalies are labelled and coloured. Note however that anomalies resulting from the generating Gaussian distribution for each of the clusters are not highlighted as such since it would require selecting an anomaly detection algorithm which may give results different to another, and that such algorithms often require appropriate but unknown parameters. In addition to the labelled scatter plot, a histogram is also provided to show the relative densities which is otherwise hidden in the plot.
@@
Figure 1: Three labelled Gaussian distributed clusters with a few additional mislabelled examples, some isolated anomalies and a small anomalous cluster.
@@
A tiny fraction of the data—$35$ examples ($0.2\%$) together with their labels—are randomly chosen from the data (excluding the small anomalous cluster). These labels act as the seeds for the proposed semi-supervised clustering algorithm and are expected to have been provided by the user after familiarising themselves and interacting with the data and domain. Note that the seeds need not cover all clusters since unrecognised clusters will simply be reported as anomalies and can be analysed in subsequent rounds. The results of the algorithm are shown in Figure 2 where three clusters are identified along with the isolated anomalies and small anomalous cluster; the latter of which can be investigated by the user in order to decide if it satsifies her criteria of being a new cluster. Note that the algorithm importantly flags the points in the outer regions—which we name fringe points—of each cluster as anomalous. Closer inspection of one of the clusters and its histogram distribution illustrates how this can be desirous since these points are strong contenders for being considered outliers compared to the main mass of the cluster (Figure 3). Thus, the proposed method carries out a clustering of data as opposed to partitioning all data points into some group. Indeed, $k$-means partitions all data into one of three clusters—including the anomalies (Figure 4). DBSCAN, on the other hand produces arguably good results here as it largely detects the three clusters and also produces anomaly labels for points that were deemed insufficient to be part of a cluster (Figure 5).
Figure 2: The proposed semi-supervised clustering algorithm successfully identifies the three underlying clusters while simultaneously detecting anomalies. Notably, the anomalous cluster is not forced into any of the labelled clusters, preserving the integrity of both the clustering and anomaly detection processes.
@@
Figure 3: A close up view of cluster 1 from Figure 2 showing the identification of not only anomalies and mislabelled data, but also fringe points that can be considered rare or unusual in a Gaussian distributed cluster of points.
@@
Figure 5: DBSCAN results using its default parameters are shown, where it largely detects the three clusters and also produces anomaly labels for points that were deemed insufficient to be part of a cluster.